32 The price elasticity of demand for a certain agricultural product is constant (over the relevant range of and equal to-2. The supply elasticity for this product is constant and equal to 3. Originally the prices) equilibrium price of this good product was unhealthy. The quantity that would be demanded at any price fell by 100%. The percent change in the long-run equilibrium consumption of this good was was $45 per unit. Then it was discovered that consumption of this
Answer #1

Solution: -60%

Explanation:

We are given that there is a fall in demand, i.e. the quantity demanded at every price falls by 100%

Supply elasticity = Percentage in quantity supplied/Percentage change in price

3 = -100 / Percentage change in price

Percentage change in price = -100 / 3 =-33.33%

Equilibrium price = (1-0.33) *45 = 30

Price elasticity = Percentage change in quantity demanded/Percentage change in price

-2 = Percentage change in quantity demanded / 30

Percentage change in quantity demanded = 2*30 = 60%